If you have a solution that is .25 orange juice and a solution that is 100% pure orange juice how much of each to get 17 liters that is 60% orange juice.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

If you have a solution that is .25 orange juice and a solution that is 100% pure orange juice how much of each to get 17 liters that is 60% orange juice.
====
.25 = 25%
Do it like this one.
----
One canned Juice drink is 15% orange juice, another is 5% orange juice. How many liters of each should be mixed together to get 10L that is 6% orange juice.
---
t = 15%
f = 5%
---
t + f = 10 (total solution)
15t + 5f = 10*6 (total OJ)
Solve for t & f

Answer by stanbon(75887)

If you have a solution that is .25 orange juice and a solution that is 100% pure orange juice how much of each to get 17 liters that is 60% orange juice.
------
Equation:
OJ + OJ = OJ
----
0.25x + 1.00(17-x) = 0.70*17
------------------
25x + 100*17 - 100x = 70*17
----
-75x = -30*17
----
x = (2/5)(17) = 6.8 liters (amt. of 25% solution needed)
----
17-x = 10.2 liters (amt. of pure OJ needed)
-----------
Cheers,
Stan H.
------------